LUCIFER Commissioning - 2008 December 10 UT

Observers: W. Seifert, N. Ageorges, M. Jutte, J. Hill (on mtn)
Telescope Operator: A. Ceranski
Software Support: T. Sargent on-call.

Summary

We were open about 45 minutes. Essentially the whole night was lost to weather.

Details

Preparation

Ran snow melting from 10 AM to 4 PM to clear the roof. Still a drift on the right shutter door, but it is stabilized. John L. knocked all the hoarfrost that he could reach off of the lightning mast and the weather mast.

Ran dome cooling for much of the previous night and day, but chamber was still at +3 degC. Started cooling the dome with the right vent door around 3PM. We closed the vent door when we opened the main shutters. The right side door is fairly clear, but the others have ice and snow.

Have been running mirror ventilation since 4PM, and mirrors are down to +1 degC for opening.

After all the work of replacing the pump on HBS motor #4, we found that last night's problem was that the motor was running backwards. You can set the wrong direction on the VFD control panel, and this must have happened over the weekend.

The larger (1/4 HP) circulating pump is installed on the oil tank, but we are still struggling with hot return oil bypassing the cold slug of oil in the back of the tank.

ECS had to be restarted in the early afternoon. We think it got confused when Bruce loaded in some new PLC code.

Open

01:26 Open Dome, T=-8.7 degC, humidity high (ice on sensors), wind is 5-8 m/sec from North, sky is clear but not photometric. We opened ~50 minutes late because of the weather conditions and mirror temperatures.

01:34 Nancy taking LUCIFER flats.

Close

01:54 Close Dome due to rising humidity and low clouds blowing over from the North side of the mountain. The good news is that the chamber is cooled off, so we can work on getting the mirrors cooled down (presently at 0 degC).

Open

03:42 Open again as humidity is starting to trend down. T=-9.3 degC. No sign of low clouds. Some blowing snow crystals at ground level.

Close

04:10 Close again as the wind is increasing (LBT: 10-15 m/sec, SMT: not updating) so there is diamond dust in the air in the dome. The only good news is that each of these open/close cycles cools the dome and mirror another 2 degC. Dome is at -4 degC, and mirrors are at -3 degC. The Polaris monitor is reporting 4 arcsec seeing (maybe influenced by wind?).

-- JohnHill - 10 Dec 2008
Topic revision: r3 - 11 Dec 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback